OpenStudy (alexandrea.lane):

HELP PLEASE! Find the geometric mean of 15 and 20

OpenStudy (alexandrea.lane):

\[\frac{ 15 }{ x }=\frac{ x }{ 20 }\]

OpenStudy (alexandrea.lane):

\[x^{2}=300\]

OpenStudy (alexandrea.lane):

\[x=\sqrt{300}\]

jhonyy9 (jhonyy9):

Find the geometric mean of 15 and 20 so do you know what mean this ?

OpenStudy (alexandrea.lane):

yes

jhonyy9 (jhonyy9):

so than - what ? what you have wrote above - is sure ? so than what i need help you ?

OpenStudy (alexandrea.lane):

I forget how to simply the square root.

OpenStudy (alexandrea.lane):

Nevermind.. I got it

jhonyy9 (jhonyy9):

sqrt300 = sqrt(3*100) = ? 100 = 10^2

jhonyy9 (jhonyy9):

so what have you get ?

OpenStudy (alexandrea.lane):

\[10\sqrt{3}\]

jhonyy9 (jhonyy9):

exactly - nice job - good luck